OpenStudy (anonymous):

2. Perform the indicated operation and, if possible, simplify -5/9 – (-2/3)

OpenStudy (anonymous):

-5/9 +2/3 = -5/9+6/9 = 1/9

OpenStudy (anonymous):

thank you